OpenStudy (anonymous):

somebody help! A= n*n matrix, det (A) = 7, find det(2A)

OpenStudy (anonymous):

det cA = c^n detA

OpenStudy (anonymous):

2^n(det(A)) =2^n(7)

OpenStudy (anonymous):

yeh i knew to use that formula...just wasn't sure if the answer was suppose to be like 2^n(7)

OpenStudy (anonymous):

or something else

OpenStudy (anonymous):

If you were given the value of n then the answer would've been different.

OpenStudy (anonymous):

multiblying A by 2 A=2n*n det(A)=14

OpenStudy (anonymous):

ah k...excellent...thanks...thought it was a trick question

OpenStudy (amistre64):

a b c d det = ad - bc 2a 2b 2c 2d det = 4ad - 4bc = 2^2(ad-bc)

OpenStudy (zarkon):

\[2^n\times 7\]